Please help solve:
There exists x in Q, such that you have 3x^2-27=0. Q= rational numbers

Answer by IWork4Dessert(57)

This problem is very fancily worded, but is really very simple. All you need to do is solve the equation.
3x^2-27=0
Move 27 to the other side...
3x^2=27
Divide both sides by 3 to get the variable alone...
x^2=9
Solve for x.
x=3
Hope this helps :)